# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

### v4

@@ 6.4. Detecting False Mappings
LLMs are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs are non-trivial mappings that the classical text preprocessing methods cannot detect.
Figure 11 shows the discovery rate of FPs using LLMs on the Largebio Track. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classical text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
@@ 6.5. Implications
LLMs were originally developed and trained for question answering (QA). OM is not only a QA task, but also a knowledge-intensive task that requires a comprehensive understanding and reasoning of domain knowledge. The studies in this section demonstrate the limitations and opportunities of integrating the text preprocessing pipeline with LLMs in syntactic OM. We highlight two pathways for this integration. One pathway is to use function calling to invoke the classical text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use. This provides stable results for text preprocessing in syntactic OM. The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction. This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs so that they can better understand the context of the task (i.e. fine-tuning LLMs used in syntactic OM).
@@ 7. Discussion
OM aims to bridge the gap by capturing similar concepts that occur in different ontologies. OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on a common practice using the text preprocessing pipeline in syntactic OM. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@
    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods. Such experimental results will benefit decision-making when selecting appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
@@
    The use of LLMs is not a one-size-fits-all solution for OM tasks. OM is a complex multi-level matching mixed with different matching techniques. LLMs may not outperform traditional methods in performing these sub-matching tasks. We suggest “rules of thumb” for syntactic OM. If two ontology entities can be matched using the intuitive text preprocessing pipeline, there is no need to leverage LLMs for the same purpose.
@@
    The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper. We refer the reader to several recent works in (He et al., 2023; Norouzi et al., 2023; Hertling and Paulheim, 2023; Qiang et al., 2023; Giglou et al., 2024; Amini et al., 2024).
  5. (5)
    This study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
@@ 8. Limitations
The references contained in the OAEI tracks are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between classes and properties. Subsumption mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings. This work has not been included in the official OAEI campaign because of its intense focus on only text preprocessing in syntactic OM.
@@ 9. Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI tracks with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. recall) and correctness (i.e. precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. recall), but matching correctness (i.e. precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. precision) and overall matching performance (i.e. F1 score). (3) We also argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs. It is an integral part of the LLMs used in OM tasks. Integrating the classical text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches. Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
@@ Acknowledgements.
The authors thank Sven Hertling for curating the Ontology Alignment Evaluation Initiative (OAEI) datasets. The authors thank Ernesto Jimenez-Ruiz for helpful advice on reproducing the OAEI Largebio Track. The authors thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v5

@@ 6.4. Detecting False Mappings
LLMs are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs are non-trivial mappings that cannot be detected by the classic text preprocessing method.
Figure 11 shows the discovery rate of FPs using LLMs on the Largebio Track. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classic text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
@@ 6.5. Implications
LLMs were originally developed and trained for question answering (QA). OM is not only a QA task, but also a knowledge-intensive task that requires a comprehensive understanding and reasoning of domain knowledge. The studies in this section demonstrate the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM. We highlight two pathways for this integration. One pathway is to use function calling to invoke the classic text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use. This provides stable results for text preprocessing in syntactic OM. The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction to reduce the number of FPs. This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and to feed the validated data back into LLMs so that they can better understand the context of the task (i.e. fine-tuning LLMs for syntactic OM).
@@ 7. Discussion
OM aims to bridge the gap by capturing similar concepts that occur in different ontologies. OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@
    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods. Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
@@
    The use of LLMs is not a one-size-fits-all solution for OM tasks. OM is a complex multi-level matching mixed with different matching techniques. LLMs may not outperform traditional methods in performing these sub-matching tasks. We suggest “rules of thumb” for syntactic OM. If two ontology entities can be matched using the intuitive text preprocessing pipeline, there is no need to leverage LLMs for the same purpose. The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper. We refer the reader to several recent works in (He et al., 2023; Norouzi et al., 2023; Hertling and Paulheim, 2023; Qiang et al., 2023; Giglou et al., 2024; Amini et al., 2024).
@@
    This study demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
@@ 8. Limitations
The references contained in the OAEI tracks are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between classes and properties. Subsumption mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings. This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
@@ 9. Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI tracks with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. recall) and correctness (i.e. precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. recall), but matching correctness (i.e. precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. precision) and overall matching performance (i.e. F1 score). (3) We argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs. It is an integral part of the LLMs used in OM tasks. Integrating the classic text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches. Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
@@ Acknowledgements.
The authors thank Sven Hertling for curating the Ontology Alignment Evaluation Initiative (OAEI) datasets. The authors thank Ernesto Jimenez-Ruiz for helpful advice on reproducing the OAEI Largebio Track. The authors thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project. According to the OAEI data policy (retrieved January 1, 2025), “OAEI results and datasets, are publicly available, but subject to a use policy similar to the one defined by NIST for TREC. These rules apply to anyone using these data.” Please find more details from the official website: https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.